# Beamforming and Resource Allocation for Delay Optimization in RIS-Assisted OFDM Systems

### v3

@@ II-B Queueing Model
In practice, packets arrive randomly at the transmitter and stored provisionally in a buffer before transmission to the receiver. Following the recommendation of the 3rd generation partnership project (3GPP), we model the packet arrival process as a Poisson process [32]. Specifically, the number of packets arriving at the buffer $k$ during time slot $t$, denoted by $\ell_{k}[t]$, follows the probability density function
@@
where $\lambda_{k}$ is the average packet arrival rate of the user $k$. The number of packets that can be theoretically transmitted from the buffer $k$ to the user $k$ in the time slot $t$ is denoted as $D_{k}[t]$ given by
@@
where $q_{k}\left[t\right]$ denotes the number of backlogged packets in the buffer $k$ at the beginning of time slot $t$. Thus, the queue in the buffer $k$ updates according to
@@ II-C Problem Formulation
The delay experienced by each packet consists of two components: queueing delay and transmission delay, both of which are affected by the subcarrier allocation indicator, the RIS reflection phase shift, and the beamforming vector at the base station. Therefore, both delay components are considered in this paper. Let $m(k)$ represent the number of packets successfully delivered to the user $k$ during the communication period with $T_{\text{sum}}$ timeslots. The delay of the $g$-th packet of the user $k$ during the considered time duration is denoted as $T_{D}^{(k)}(g)$ for all $g\in\{1,2,...,m(k)\}$, which accounts for both queueing and transmission delays.
In this paper, our objective is to minimize the average delay by jointly optimizing the subcarrier allocation shift $\alpha_{n,k}[t]$, beamforming vectors at the BS $\mathbf{w}_{n}[t]$, and RIS reflection phase shifts $\theta_{m}[t]$. Specifically, the problem is formulated as
@@
where (11a) describes the binary constraints of the subcarrier allocation indicaters, (11b) demonstrates that each subcarrier can be allocated to at most one user, (11c) ensures that the total power consumption of the BS does not exceed its maximum budget $P_{\max}$, and (11d) limits each RIS element’s reflection phase shift to its allowable range.
III Deep Reinforcement Learning Based Algorithm
This section introduces a DRL-based framework for addressing the problem in (11), highlighting the core design principles, the corresponding training methodology and the computational complexity of proposed algorithm.
@@ III-A Overall Framework
The optimization problem presented in equation (11) necessitates determining the RIS phase shifts and resource allocation for each time slot, where decisions made in the current slot directly influence the number of packets that can be delivered, thereby affecting the queue length in the subsequent slot. Therefore, the problem (11) is a sequential decision making issue, which can be typically characterized by MDP. To this end, a hybrid RL approach is adopted, wherein a central controller collects information regarding the channel state information and the states of all buffers. The central controller, deployed at the BS, functions as an agent responsible for controlling the RIS and subcarrier allocation. It explores the unknown communication environment to acquire experience, which is subsequently utilized to guide policy design and optimize the RIS reflection phase shift and subcarrier allocation strategies based on observations of the environment state. Specifically, the PPO-$\Theta$ is utilized to predict the RIS reflection phase shift, while PPO-$N$ is applied to determine the subcarrier allocation indicaters. The detailed algorithm framework is illustrated in Fig. 2.
@@
4:     Collect $\bar{\mathbf{h}}^{(d)}_{n,k}[t]$, $\bar{\mathbf{h}}^{(r)}_{n,k}[t]$, $q_{k}\left[t\right]$, $R_{k}[t-1]$ and $\ell_{k}\left[t\right]$.
@@
10:        Calculate ${\mathbf{w}_{n}[t]}$ according Section III-D.
@@ III-B PPO-$\Theta$ Design
The PPO-$\Theta$ model is employed to optimize the RIS reflection phase shift. Given the continuous nature of RIS phase shifts, the PPO algorithm is utilized to train an agent comprising an actor network for action selection and a critic network for state value estimation. Specifically, the actor network determines the RIS reflection phase shift, while the critic network evaluates the effectiveness of the selected policy. In the following, we introduce the state, action, and reward, which are the three fundamental elements for reinforcement learning algorithm design.
In wireless communication, channel state information significantly impacts communication performance, making it essential for RIS reflection phase shift design and subcarrier allocation. Moreover, the buffer state information, i.e., the number of queueing packets of each user, and factors that may influcence the buffer state, including the number of current arrival packets, need to be observed every slot to make sequential delay-aware decision. Therefore, the environment state $s_{t}^{(\Theta)}$ is defined as the channel state information at the time slot $t$, including the direct link channel $\bar{\mathbf{h}}^{(d)}_{n,k}[t]$ and cascaded reflection channel $\bar{\mathbf{h}}^{(r)}_{n,k}[t]$, the number of packets backlogged in each user’s buffer $q_{k}\left[t\right]$, and the number of current arrival packets $\ell_{k}\left[t\right]$.
In reinforcement learning, an action refers to the decision or operation made by an agent in response to a given state of the environment. In the proposed framework, the action of PPO-$\Theta$, denoted as $a_{t}^{(\Theta)}$, is defined as the RIS reflection phase shift vector $\Theta[t]=\{\theta_{1}[t],\theta_{2}[t],\cdots,\theta_{M}[t]\},$.
@@
b) Critic network design: The critic network adopts the same input structure and feature extraction architecture as those of the actor network. However, it differs in its final layer, which is tailored to output a scalar value estimating the state value $V_{\phi^{(\Theta)}}(s_{t}^{(\Theta)})\in\mathbb{R}$.
(a) Actor and critic network architectures of PPO-$\Theta$. While the figure illustrates the structurally similar components of the actor and critic networks in a unified manner for clarity, the two networks are implemented with separate parameters.
@@
$\mathcal{J}(\theta^{(\Theta)})=\;\mathbb{E}_{t}\Big{[}\min\big{(}$ $\bar{r}_{t}^{(\Theta)}(\theta^{(\Theta)})\hat{A}_{t}^{(\Theta)},$ (13)
$\text{clip}(\bar{r}_{t}^{(\Theta)}(\theta^{(\Theta)}),1-\epsilon^{(\Theta)},1+\epsilon^{(\Theta)})\hat{A}_{t}^{(\Theta)}\big{)}\Big{]}$
where $\bar{r}_{t}^{(\Theta)}(\theta^{(\Theta)})=\frac{\pi_{\theta^{(\Theta)}}(a_{t}^{(\Theta)}|s_{t}^{(\Theta)})}{\pi_{\theta^{(\Theta)}_{\text{old}}}(a_{t}^{(\Theta)}|s_{t}^{(\Theta)})}$ denotes the probability ratio between the new and old policies, $\epsilon^{(\Theta)}$ is a clipping factor used in PPO to stabilize learning by preventing excessively large changes of policy update, and $\hat{A}_{t}^{(\Theta)}$ is the generalized advantage estimation (GAE). The GAE can be computed by
@@
where $\gamma$ is the discount factor, $\lambda$ is the GAE parameter, and the temporal difference (TD) error $\delta_{t}^{(\Theta)}$ is defined by
$\delta_{t}^{(\Theta)}=r_{t}+\gamma V_{\phi^{(\Theta)}}(s_{t+1}^{(\Theta)})-V_{\phi^{(\Theta)}}(s_{t}^{(\Theta)}).$ (15)
@@
$\mathcal{L}(\phi^{(\Theta)})=\mathbb{E}_{t}\left[\left(r_{t}+\gamma V_{\phi^{(\Theta)}}(s_{t+1}^{(\Theta)})-V_{\phi^{(\Theta)}}(s_{t}^{(\Theta)})\right)^{2}\right],$ (16)
where $V_{\phi^{(\Theta)}}(s_{t+1}^{(\Theta)})$ and $V_{\phi^{(\Theta)}}(s_{t}^{(\Theta)})$ represent the value estimations of the next and current states, respectively.
@@ III-C PPO-$N$ Design
The global state $s_{t}^{(N)}$ used by the centralized critic includes the complete equivalent channel state information $\mathbf{h}_{n,k}^{\text{eff}}$ for all users and subcarriers, along with the global buffer state $q_{k}[t]$ and packet arrivals $\ell_{k}[t]$, enabling joint policy evaluation across agents. To reduce input dimensionality, $\mathbf{h}_{n,k}^{\text{eff}}$ is defined as $\mathbf{\bar{h}}^{(d)}_{n,k}+\mathbf{\varphi}^{T}\mathbf{\bar{h}}^{(r)}_{n,k}$, where $\mathbf{\varphi}[t]$ is the RIS reflection phase shifts vector generated by PPO-$\Theta$. The local observation $o_{t}^{(n)}$ of agent $n$ comprises the equivalent channel state from the BS to users on subcarrier $n$, buffer length $q_{k}[t]$, and packet arrivals $\ell_{k}[t]$.
Since the objective is to minimize average delay by jointly optimizing RIS phase shifts and subcarrier allocation, PPO-$N$ and PPO-$\Theta$ share the reward function defined in (12). Since $N$ agents are used to allocate $N$ subcarriers, and each subcarrier has $K$ possible allocation options, the action space $a_{t}^{(n)}$ of each agent is defined as $\{1,2,\dots,K\}$.
@@
$\mathcal{J}(\theta^{(n)})=$ $\;\mathbb{E}\Big{[}\min\big{(}\bar{r}_{t}(\theta^{(n)})\hat{A}_{t}^{(n)},$ (17)
$\quad\text{clip}(\bar{r}_{t}^{(n)}(\theta^{(n)}),1-\epsilon^{(n)},1+\epsilon^{(n)})\hat{A}_{t}^{(n)}\big{)}\Big{]}$
@@
$\bar{r}_{t}^{(n)}(\theta^{(n)})=\frac{\pi_{\theta^{(n)}}(a_{t}^{(n)}|o_{t}^{(n)})}{\pi_{\theta_{\text{old}}^{(n)}}(a_{t}^{(n)}|o_{t}^{(n)})}.$ (18)
$\hat{A}_{t}^{(n)}$ is the GAE for agent $n$, which is computed as
@@
$\delta_{t}^{(n)}=r_{t}+\gamma V_{\phi^{(N)}}(s_{t+1}^{(N)})-V_{\phi^{(N)}}(s_{t}^{(N)}).$ (20)
@@
$\mathcal{L}(\phi^{(N)})=\mathbb{E}_{t}\left[\left(r_{t}^{(n)}+V_{\phi^{(N)}}(s_{t+1}^{(N)})-V_{\phi^{(N)}}(s_{t}^{(N)})\right)^{2}\right],$ (21)
where $V_{\phi^{(N)}}(s_{t+1}^{(N)})$ and $V_{\phi^{(N)}}(s_{t}^{(N)})$ represent the value estimations of the next and current global states, respectively, and $r_{t}^{(n)}$ denotes the reward obtained by agent $n$ from subcarrier allocation decisions.
@@ III-D Active Beamforming Design
With the RIS reflection shift and subcarrier allocation indicater produced by PPO-$\Theta$ and PPO-$N$, the BS beamforming vectors ${\mathbf{w}_{n}[t]}$ for subcarrier $n$ at the time slot $t$ can be derived as maximum ratio transmission (MRT) in conjunction with the water-filling algorithm, i.e.,
@@
where $\bar{\mathbf{w}}_{n,q}$ represents the optimal unit-norm beamforming vector and $p_{n}[t]$ refers to the transmission power on subcarrier $n$ at the time slot $t$. In particular, $\bar{\mathbf{w}}_{n,q}$ can be expressed as,
@@
The $p_{n}[t]$ can be expressed by the following equation,
@@
where $c_{n}[t]=\frac{\|(\bar{\mathbf{h}}^{(d)}_{n,k}[t]+\boldsymbol{\varphi}[t]^{T}\bar{\mathbf{h}}^{(r)}_{n,k}[t])\|^{2}}{\sigma^{2}}$, and $\tau[t]$ is the Lagrange multiplier satisfying,
@@ III-E Training
Since the RIS phase has not been optimized in the early stages of training, the system experiences low data rates and overall instability. As a result, the reward value decreases as the time step increases within an episode, making training difficult. To address this, this paper proposes a novel transfer reinforcement learning strategy. Specifically, in the initial training phase, the reward is set to $\min R_{k}[t]$, while the training process still follows Algorithm 1. The trained PPO-$\Theta$ model parameters are then saved and reloaded into Algorithm 1 as the new PPO-$\Theta$ model initialization parameters. The reward is then switched to (12), and training continues until convergence through continuous iterative optimization.

### v4

@@ II-B Queueing Model
In practice, packets arrive randomly at the transmitter and stored provisionally in a buffer before transmission to the receiver. Following the recommendation of the 3rd generation partnership project (3GPP), we model the packet arrival process as a Poisson process [35]. Specifically, the number of packets arriving at buffer $k$ during time slot $t$, denoted by $\ell_{k}[t]$, follows the probability density function
@@
where $\lambda_{k}$ is the average packet arrival rate of user $k$. The number of packets that can be theoretically transmitted from the buffer $k$ to user $k$ in time slot $t$ is denoted as $D_{k}[t]$ given by
@@
where $q_{k}\left[t\right]$ denotes the number of backlogged packets in buffer $k$ at the beginning of time slot $t$. Thus, the queue in the buffer $k$ is updated according to
@@ II-C Problem Formulation
The delay experienced by each packet consists of two components: queueing delay and transmission delay, both of which are affected by the subcarrier allocation indicator, RIS reflection phase shifts, and the beamforming vector at the BS. Therefore, both delay components are considered in this paper. Let $m(k)$ represent the number of packets successfully delivered to user $k$ during the communication period with $T_{\text{sum}}$ time slots. The delay of the $g$-th packet of user $k$ during the considered time duration is denoted as $T_{D}^{(k)}(g)$ for all $g\in\{1,2,...,m(k)\}$, which accounts for both queueing and transmission delays.
In this paper, our objective is to minimize the average delay by jointly optimizing the subcarrier allocation indicator, $\alpha_{n,k}[t]$, beamforming vectors at the BS, $\mathbf{w}_{n}[t]$, and RIS reflection phase shifts, $\theta_{m}[t]$. Specifically, the problem is formulated as
@@
where (11a) describes the binary constraints of the subcarrier allocation indicaters, (11b) demonstrates that each subcarrier can be allocated to at most one user, (11c) ensures that the total power consumption of the BS does not exceed its maximum budget $P_{\max}$, while (11d) limits each RIS element’s reflection phase shift to its allowable range.
III DRL-Based Algorithm
This section introduces a DRL-based framework for addressing the problem in (11), highlighting the core design principles, corresponding training methodology and the computational complexity of the proposed algorithm.
@@ III-A Overall Framework
The optimization problem presented in equation (11) necessitates determining the RIS phase shifts and resource allocation for each time slot, where decisions made in the current slot would directly influence the number of packets that can be delivered, thereby affecting the queue length in the subsequent slot. Therefore, problem (11) is a sequential decision making issue, which can be typically characterized by MDP. To this end, a hybrid RL approach is adopted, wherein a central controller collects information regarding the CSI and the states of all buffers. The central controller, deployed at the BS, functions as an agent responsible for controlling the RIS and subcarrier allocation. It explores the unknown communication environment to acquire experience, which is subsequently utilized to guide the policy design and optimize the RIS reflection phase shifts and subcarrier allocation strategies based on observations of the environment state. Specifically, the PPO-$\Theta$ is utilized to predict the RIS reflection phase shifts, while PPO-$N$ is applied to determine the subcarrier allocation indicators. The detailed algorithm framework is illustrated in Fig. 2.
@@
4:     Collect $\bar{\mathbf{h}}^{(d)}_{n,k}[t]$, $\bar{\mathbf{H}}^{(r)}_{n,k}[t]$, $q_{k}\left[t\right]$, and $\ell_{k}\left[t\right]$.
@@
10:        Calculate ${\mathbf{w}_{n}[t]}$ according to Section III-D.
@@ III-B PPO-$\Theta$ Design
The PPO-$\Theta$ model is employed to optimize the RIS reflection phase shifts. Given the continuous nature of RIS phase shifts, the PPO algorithm is utilized to train an agent comprising an actor network for action selection and a critic network for state value estimation. Specifically, the actor network determines the RIS reflection phase shifts, while the critic network evaluates the effectiveness of the selected policy. In the following, we introduce the state, action, and reward, which are the three fundamental elements for algorithm design.
In wireless communication, CSI significantly impacts the communication performance, making it essential for RIS reflection phase shift design and subcarrier allocation. Moreover, the buffer state information, i.e., the number of queueing packets of each user, and factors that may influcence the buffer state, including the number of current arrival packets, need to be observed in every time slot to make sequential delay-aware decisions. Therefore, the environment state $s_{t}^{(\Theta)}$ is defined as the CSI at time slot $t$, including the direct link channel $\bar{\mathbf{h}}^{(d)}_{n,k}[t]$ and cascaded reflection channel $\bar{\mathbf{H}}^{(r)}_{n,k}[t]$, the number of packets backlogged in each user’s buffer $q_{k}\left[t\right]$, and the number of current arrival packets $\ell_{k}\left[t\right]$.
In RL, an action refers to the decision or operation made by an agent in response to a given state of the environment. In the proposed framework, the action of PPO-$\Theta$, denoted as $a_{t}^{(\Theta)}$, is defined as the RIS reflection phase shift vector $\boldsymbol{\theta}[t]=\{\theta_{1}[t],\theta_{2}[t],\cdots,\theta_{M}[t]\}$.
@@
b) Critic network design: The critic network adopts the same input structure and feature extraction architecture as those of the actor network. However, it differs in its final layer, which is tailored to output a scalar value estimating the state value $V_{\phi^{(\Theta)}}\left(s_{t}^{(\Theta)}\right)\in\mathbb{R}$.
(a) Actor and critic network architectures of PPO-$\Theta$. While the figure illustrates the structurally similar components of the actor and critic networks in a unified manner for clarity, the two networks are implemented with separate parameters
@@
[…]
$\text{clip}(\bar{r}_{t}^{(\Theta)}(\theta^{(\Theta)}),1-\epsilon^{(\Theta)},1+\epsilon^{(\Theta)})\hat{A}_{t}^{(\Theta)}\bigg{)}\Bigg{\}},$
where $\bar{r}_{t}^{(\Theta)}(\theta^{(\Theta)})=\frac{\pi_{\theta^{\left(\Theta\right)}}\left(a_{t}^{(\Theta)}|s_{t}^{(\Theta)}\right)}{\pi_{\theta^{(\Theta)}_{\text{old}}}\left(a_{t}^{(\Theta)}|s_{t}^{(\Theta)}\right)}$ denotes the probability ratio between the new and old policies, $\epsilon^{(\Theta)}$ is a clipping factor used in PPO to stabilize learning by preventing excessively large changes of policy update, and $\hat{A}_{t}^{(\Theta)}$ is the generalized advantage estimation (GAE). The GAE can be computed by
@@
where $\gamma$ is the discount factor, $\lambda$ is the GAE parameter, while the temporal difference (TD) error $\delta_{t}^{(\Theta)}$ is defined by
$\delta_{t}^{(\Theta)}=r_{t}+\gamma V_{\phi^{(\Theta)}}\left(s_{t+1}^{(\Theta)}\right)-V_{\phi^{(\Theta)}}\left(s_{t}^{(\Theta)}\right).$ (15)
@@
$\mathcal{L}(\phi^{(\Theta)})=\mathbb{E}_{t}\left\{\bigg{(}r_{t}+\gamma V_{\phi^{(\Theta)}}\left(s_{t+1}^{(\Theta)}\right)-V_{\phi^{(\Theta)}}\left(s_{t}^{(\Theta)}\right)\bigg{)}^{2}\right\},$ (16)
where $V_{\phi^{(\Theta)}}\left(s_{t+1}^{(\Theta)}\right)$ and $V_{\phi^{(\Theta)}}\left(s_{t}^{(\Theta)}\right)$ represent the value estimations of the next and current states, respectively.
@@ III-C PPO-$N$ Design
The global state $s_{t}^{(N)}$ used by the centralized critic includes the complete equivalent channel state information $\mathbf{h}_{n,k}^{\text{eff}}[t]$ for all users and subcarriers, along with the global buffer state $q_{k}[t]$ and packet arrivals $\ell_{k}[t]$, enabling joint policy evaluation across agents. To reduce the input dimensionality, $\mathbf{h}_{n,k}^{\text{eff}}[t]$ is defined as $\mathbf{\bar{h}}^{(d)}_{n,k}[t]+\boldsymbol{\mathbf{\varphi}}[t]^{T}\mathbf{\bar{H}}^{(r)}_{n,k}[t]$. The local observation $o_{t}^{(n)}$ of agent $n$ comprises the equivalent channel state from the BS to users on subcarrier $n$, buffer length $q_{k}[t]$, and packet arrivals $\ell_{k}[t]$.
Since the objective is to minimize the average delay by jointly optimizing the RIS phase shifts and subcarrier allocation, PPO-$N$ and PPO-$\Theta$ share the reward function defined in (12). Since $N$ agents are used to allocate $N$ subcarriers, and each subcarrier has $K$ possible allocation options, the action space $a_{t}^{(n)}$ of each agent is defined as $\{1,2,\dots,K\}$.
@@
$\mathcal{J}(\theta^{(n)})=$ $\;\mathbb{E}\Bigg{\{}\min\bigg{(}\bar{r}_{t}\left(\theta^{(n)}\right)\hat{A}_{t}^{(n)},$ (17)
$\quad\text{clip}\left[\bar{r}_{t}^{(n)}\left(\theta^{(n)}\right),1-\epsilon^{(n)},1+\epsilon^{(n)}\right]\hat{A}_{t}^{(n)}\bigg{)}\Bigg{\}},$
@@
$\bar{r}_{t}^{(n)}(\theta^{(n)})=\frac{\pi_{\theta^{(n)}}\left(a_{t}^{(n)}|o_{t}^{(n)}\right)}{\pi_{\theta_{\text{old}}^{(n)}}\left(a_{t}^{(n)}|o_{t}^{(n)}\right)}.$ (18)
Note that $\hat{A}_{t}^{(n)}$ is the GAE for agent $n$, which is computed as
@@
$\delta_{t}^{(n)}=r_{t}+\gamma V_{\phi^{(N)}}\left(s_{t+1}^{(N)}\right)-V_{\phi^{(N)}}\left(s_{t}^{(N)}\right).$ (20)
@@
$\mathcal{L}(\phi^{(N)})=\mathbb{E}_{t}\bigg{\{}\bigg{(}r_{t}^{(n)}+V_{\phi^{(N)}}\left(s_{t+1}^{(N)}\right)-V_{\phi^{(N)}}\left(s_{t}^{(N)}\right)\bigg{)}^{2}\bigg{\}},$ (21)
where $V_{\phi^{(N)}}\left(s_{t+1}^{(N)}\right)$ and $V_{\phi^{(N)}}\left(s_{t}^{(N)}\right)$ represent the value estimations of the next and current global states, respectively, and $r_{t}^{(n)}$ denotes the reward obtained by agent $n$ from subcarrier allocation decisions.
@@ III-D Active Beamforming Design
With the RIS reflection shift and subcarrier allocation indicators produced by PPO-$\Theta$ and PPO-$N$, the BS beamforming vectors ${\mathbf{w}_{n}[t]}$ for subcarrier $n$ at the time slot $t$ can be designed by leveraging the maximum ratio transmission (MRT) principle in conjunction with the water-filling algorithm, i.e.,
@@
where $\bar{\mathbf{w}}_{n,q}$ represents the optimal unit-norm beamforming vector and $p_{n}[t]$ refers to the transmission power on subcarrier $n$ at time slot $t$. In particular, $\bar{\mathbf{w}}_{n,q}$ can be expressed as,
@@
The term $p_{n}[t]$ can be expressed by the following equation,
@@
where $c_{n}[t]=\frac{\|(\bar{\mathbf{h}}^{(d)}_{n,k}[t]+\boldsymbol{\varphi}[t]^{T}\bar{\mathbf{H}}^{(r)}_{n,k}[t])\|^{2}}{\sigma^{2}}$, while $\tau[t]$ is the Lagrange multiplier satisfying,
@@ III-E Training
Since the RIS phases have not been optimized in the early stages of training, the system experiences low data rates and overall instability. As a result, the reward value decreases as the time step increases within an episode, making training difficult. To address this, this paper proposes a novel transfer reinforcement learning strategy. Specifically, in the initial training phase, the reward is set to $\min R_{k}[t]$, while the training process still follows Algorithm 1. The trained PPO-$\Theta$ model parameters are then saved and reloaded into Algorithm 1 as the new PPO-$\Theta$ model initialization parameters. The reward is then switched to (12), and training continues until convergence through continuous iterative optimization.